APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday June 6, 2004 (MJD 53163) ---=== OBSERVING TEAM ===--- Swing: Atsuko Nitta Night: Pete Newman Dan Long Visitors: A cohort of presidents, deans and others from NMSU toured with Bruce Gillespie. ---=== OBSERVING PLAN ===--- 12 more dark runs, 12 more shake downs to go. But just one shake for tonight. ---=== OBSERVING SUMMARY ===--- Good pointing model obtained. Both PRs on the shakedown list tested and closed. No science data due to clouds, moon and dust. The software mask for the infrared sky camera may need adjustment, and the DIMM is shutdown. ---=== OBSERVING LOG ===--- Afternoon ========= engCam mounted. I set the counter weights at 90. The past dark run, it seems like we are favoring lower counter weights for all inst. compared to what is on the box. We were able to take an image at the first try. yupeeeee. Everything else checked out fine. Night ===== Clouds from afternoon storms started to clear around sunset, so we opened as normal. We noted that the software mask for the infrared sky camera is not aligned with the tripod in the undifferentiated images (though this makes no difference in the differential images). During opening, we tried to get the DIMM running before recalling that there is a problem with the software driver for its mouse. We shutdown the computer during this (intending a restart, not a shutdown) and left it so. First order of business was a pointing model. A rotation test showed the engineering camera to be centered on the boresight to within a pixel. A 72-point data set was obtained using the GRID_DEFAULT_72PT.DAT grid. The mean temperature during the observations was 16.5+-1.5 C. The fit showed a significant change in the PZZ1 parameter. * inmod tdat:telmod SYS$SYSDEVICE:[TCC.PTERR]PTERR_20040521.DAT;1 T 41 1.5500 IE -239.9177 IA -251.7573 AN -6.0812 AW +3.1767 HWCA2 +0.6975 HWSA2 -0.2803 HNCA1 +1.0513 HNSA1 -0.0239 CA +10.9025 PZZ1 -146.7684 ACEC +0.4405 ACES +1.4248 TX -1.0862 END * getdat pterr_20040607.dat Skipping point too near zenith: 208.99 31.87 345.509909 89.034433 0.009915 -0.064642 -0.042525 0.001268 4593297458.30 Read 70 data points * fix * fit coeff change value sigma 1 IE -239.92 2 IA -251.76 3 AN -6.08 4 AW +3.18 5 HWCA2 +0.70 6 HWSA2 -0.28 7 HNCA1 +1.05 8 HNSA1 -0.02 9 CA +10.90 10 PZZ1 -146.77 11 ACEC +0.44 12 ACES +1.42 13 TX -1.09 Sky RMS = 10.08 Popn SD = 10.08 * call t4014 * use * fit coeff change value sigma 1 IE +7.097 -232.82 0.911 2 IA -7.902 -259.66 0.977 3 AN +0.767 -5.31 0.293 4 AW -0.974 +2.20 0.296 5 HWCA2 -0.868 -0.17 0.388 6 HWSA2 -0.351 -0.63 0.388 7 HNCA1 -0.494 +0.56 0.438 8 HNSA1 +0.005 -0.02 0.440 9 CA -3.997 +6.91 0.727 10 PZZ1 +7.012 -139.76 3.010 11 ACEC +0.240 +0.68 0.542 12 ACES -3.600 -2.17 0.536 13 TX -0.372 -1.46 1.428 Sky RMS = 1.60 Popn SD = 1.77 * outmod tdat:telmod A 25-point check was made which was fit by the new model with 2.0 arcsec rms error. We then proceeded to spectroscopy for science and to test changes to SOP. 06:29Z Plate 1821, cartridge 3. Main survey. Well centered guide stars on gotoField with cleared offsets confirms the pointing model is good. Seeing degraded as thw wind swung to the southeast around moonrise. No useful data obtained due to increasing cloud and moonlight. Rising dust counts and humidity dictated closure at 07:43Z. During the spectroscopy, we tested and passed the changes in PR6044 (SOP guider errors). We also confirmed that the fix to MOP under PR4195 (no dewar pressure errors) works as expected. Both PRs were closed, and IOP v3_138_0 was declared current. We gave up an hour before morning twilight. endNight ran successfully onto cloudy-night tapes. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 4658 23:20Z 23:58Z 100 O -45.01 -35.46 77 ignore doghouse ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): y QA Procedures Done (y/n): n UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 27062, plate -9999 ------- 23:24 27062 0.0 bias ----- sequence 27063, plate 1615 ------- 23:31 27063 10.0 flat 23:33 27064 2.0 arc ----- sequence 27065, plate -9999 ------- 02:09 27065 0.0 bias ----- sequence 27066, plate 1821 ------- 06:36 27066 10.0 flat 06:39 27067 2.0 arc 07:02 27068 1200.1 target Ignore 07:30 27069 1500.1 target Ignore ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 06:55Z 1 1821 -36.17 -0.0001 58.70 0.0016 149.95 -0.0216 1.000190 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- No tapes to ship or store. ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 06:55Z 1 1821 1.00019 -1876 -518 -120 -36 58.8 15.3 18 138 BG38 2.0 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:12Z 65 24 11 84 (E) 1441 - Partly cloudy 23:42Z 64 25 10 73 (ENE) 1211 - " 00:12Z 65 23 12 81 (E) 945 - " 00:43Z 65 25 6 88 (E) 1262 - " 01:15Z 63 24 9 90 (E) 1050 - " 01:47Z 62 23 7 107 (ESE) 1157 - " 02:17Z 63 30 8 134 (SE) 1880 - " 03:01Z 63 32 7 121 (ESE) 2179 - " 03:34Z 63 32 7 126 (SE) 2359 - " 04:10Z 64 28 3 184 (S) 2136 - " 04:42Z 61 28 8 127 (SE) 1928 - " 05:15Z 61 32 12 140 (SE) 2308 - " 05:47Z 61 32 8 149 (SSE) 2543 - " 06:19Z 62 34 16 139 (SE) 2600 - " 06:51Z 59 37 20 138 (SE) 2624 - " 07:23Z 59 38 18 134 (SE) 3055 - " 07:56Z 59 35 20 137 (SE) 3177 - " 08:28Z 59 33 21 150 (SSE) 2693 - " 09:00Z 58 33 17 147 (SSE) 2895 - " 09:33Z 60 28 17 165 (SSE) 2981 - ---=== TELESCOPE STATUS ===--- 02:00Z Fans on, doors open 02:20Z Enclosure off No interlocks bypassed. Status at 10:23Z: Telescope stowed at: 30 deg altitude Instrument mounted: Cartridge 1 Counterweights at: 270 Autofill systems: On 180L LN2 dewar scales: Spectro 188 lb, 23 psi Imager 188 lb, 23 psi (last fill about 04Z/22MDT) (Honest: the weights and pressures were the same!) ---=== SOFTWARE USED ===--- IOP/SOP: v3_138_0 (test, declared current) Watcher: v2_29_0 MCP: v5_25_0 TPM: tpm_v2_42_0 AstroDa: v14_47 TCC: TCC 2.7.2 November 11 2003 sdssProcedures: v1_90 SoS: v4_10_7 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.5440 0.5390 Axial B 0.6570 0.6390 Axial C 1.0330 1.0300 Trans D -8.4500 -8.4550 Lateral E 10.8839 10.8712 Lateral F 10.8458 10.8331 GALILS Commanded: 4600. -2850. 3500. -9250. 18400. 17850. Actual: 4600. -2850. 3500. -9250. 18400. 17850. SETMIR VALUES PriDesOrient: 0.00 0.00 15.20 829.90 1.60 PriOrient: 0.00 0.00 14.91 830.49 3.19 SECONDARY: ---------- Focus: 0.00 Air Temp.: 17.0 Alt.: 29.999997 MIGS TONIGHT NOMINAL Axial A 1.5090 1.5090 Axial B 1.0940 1.0960 Axial C 1.1080 1.1100 Trans D 0.0170 -0.0020 GALILS Commanded: 1628638. 1548178. 1574547. -3900. -8750. Actual: 1628650. 1548200. 1574500. -3900. -8750. SETMIR VALUES SecDesOrient: 1257.00 0.00 -25.00 0.00 146.27 SecOrient: 1257.01 -0.02 -25.00 0.26 146.57 ---=== PROBLEMS IN DETAIL ===---